The sum of the sqaures of two consecutive odd numbers is 202, what are the numbers? Show equation
Answer by mananth(16949)

let the numbers be x & x+2
x^2 +(x+2)^2 =202
x^2+x^2+4x+4=202
2x^2+4x-198=0
/2
x^2+2x-99=0
x^2+11x-9x-99=0
x(x+11)-9(x+11)=0
(x+11)(x-9)=0
the numbers are 9,11
-11,-9
...
m.ananth@hotmail.ca